OpenStudy (anonymous):

HELP!!! What is the asymptote for the above function y=2x^2/3x^2-16???

OpenStudy (anonymous):

\[3x^2-16=0\] vertical:\[x=\pm4/3*\sqrt{3} \] horizontal: \[\lim_{x \rightarrow \infty} 2x^2/(3x^2-16)=2/3\]